Two dice are rolled. What is the probability of getting a sum that equals 10?

Found 2 solutions by math_tutor2020, ikleyn:
Answer by math_tutor2020(3838)


In general,
Event space = set of desired outcomes
Sample space = set of all possible outcomes

For this particular problem:
Event space = set of dice rolls that sum to 10
Sample space = set of all possible dice rolls

There are 3 items in the event space
4+6 = 10
5+5 = 10
6+4 = 10

There are 6*6 = 36 items in the sample space.

Divide the two values

3/36 = 1/12 is the probability of getting a sum of 10.

Answer by ikleyn(53938)

.

On rolling dice problems see the lesson
    - Rolling a pair of fair dice
in this site.  You will find there many similar solved problems.


Read it and learn the subject from there.  After that,  you will be able easy solve such problems
ON  YOUR  OWN,  which is your  MAJOR  GOAL  visiting this forum  (I believe).


Also,  you have this free of charge online textbook in ALGEBRA-II in this site
    - ALGEBRA-II - YOUR ONLINE TEXTBOOK.

The referred lesson is the part of this online textbook under the topic  "Solved problems on Probability".


Save the link to this textbook together with its description

Free of charge online textbook in ALGEBRA-II
https://www.algebra.com/algebra/homework/complex/ALGEBRA-II-YOUR-ONLINE-TEXTBOOK.lesson

into your archive and use when it is needed.